OpenStudy (anonymous):

The population of a city is 83,945 and grows continuously at a rate of 1.3% each year. What is the approximate population of the city in 21 years? 110,295 105,626 110,279 106,862

OpenStudy (anonymous):

Someone please help!!!

OpenStudy (anonymous):

Note: I have no idea what to do

jimthompson5910 (jim_thompson5910):

use the formula A = Pe^(r*t) where A = population in t years P = initial population r = growth rate t = time in years

OpenStudy (anonymous):

I dont know what the "e" is

jimthompson5910 (jim_thompson5910):

e is the number 2.718... (it's like pi) and most calculators support it

OpenStudy (anonymous):

Oh foreal?

jimthompson5910 (jim_thompson5910):

if you were to type in e, your calculator should know what to do with it

jimthompson5910 (jim_thompson5910):

yep definitely real

OpenStudy (anonymous):

what if im using my iphone?

jimthompson5910 (jim_thompson5910):

then use e = 2.71828

OpenStudy (anonymous):

ok, ill try

jimthompson5910 (jim_thompson5910):

so you will go from A = Pe^(r*t) to A = 83945*2.71828^(0.013*21) and that gives you A = ??

OpenStudy (anonymous):

I know Im wrong

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

Its embarrassing... 29.03

jimthompson5910 (jim_thompson5910):

yeah way too small

jimthompson5910 (jim_thompson5910):

but that's ok, try it again

OpenStudy (anonymous):

ok, 1 minute

jimthompson5910 (jim_thompson5910):

k

OpenStudy (anonymous):

110,295?

jimthompson5910 (jim_thompson5910):

nailed it

OpenStudy (anonymous):

Thanks man!!!! :D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

If i ever need help again, can i ask you?

jimthompson5910 (jim_thompson5910):

sure, but don't hesitate to give it a try and show me what you got

OpenStudy (anonymous):

of course! catch you later man

jimthompson5910 (jim_thompson5910):

cya